An earthquake registered 6.9 on the Richter scale. It produced 3.5 x 10^13 J of energy. How much energy did the Alaska earthquake, 8.2 on the Richter scale produce?
Answer by CharlesG2(834)

An earthquake registered 6.9 on the Richter scale. It produced 3.5 x 10^13 J of energy. How much energy did the Alaska earthquake, 8.2 on the Richter scale produce?

8.2 - 6.9 = 1.3 increase in magnitude
ground movement increase = 10^(increase in magnitude)
ground movement increase = 10^1.3 = 19.952623 to 6 places
increase in energy = (ground movement increase)^1.5
increase in energy = 19.952623^1.5
increase in energy = 89.125094 to 6 places
89.125094 * 3.5 x 10^13 J = 311.937829 x 10^13 J
311.937829 x 10^13 J = approximately 3.1 x 10^15 J of energy
(using formulas off http://www.backbenchmedia.com/richter-scale-equation-earthquakes/)